# Learning Fairer Representations with FairVIC

### v2

@@ Abstract
Mitigating bias in automated decision-making systems, particularly in deep learning models, is a critical challenge due to nuanced definitions of fairness, dataset-specific biases, and the inherent trade-off between fairness and accuracy. To address these issues, we introduce FairVIC, an innovative approach that enhances fairness in neural networks by integrating variance, invariance, and covariance terms into the loss function during training. Unlike methods that rely on predefined fairness criteria, FairVIC abstracts fairness concepts to minimise dependency on protected characteristics. We evaluate FairVIC against comparable bias mitigation techniques on benchmark datasets, considering both group and individual fairness, and conduct an ablation study on the accuracy-fairness trade-off. FairVIC demonstrates significant improvements ($\approx 70\%$) in fairness across all tested metrics without compromising accuracy, thus offering a robust, generalisable solution for fair deep learning across diverse tasks and datasets.
Bias Mitigation, Fairness, Deep Learning
@@ 1 Introduction
With the ever-increasing utilisation of Artificial Intelligence (AI) in everyday applications, neural networks have emerged as pivotal tools for automated decision making systems in sectors such as healthcare (Esteva et al., 2017), finance (Dixon et al., 2017), and recruitment (Vardarlier and Zafer, 2020). However, bias in the data– stemming from historical inequalities, imbalanced distributions, or flawed feature representations–are often learned by these models, posing significant challenges to fairness. Such bias can lead to adverse decisions affecting real lives. For instance, several studies have shown how bias in facial recognition technologies disproportionately misidentifies individuals of certain ethnic backgrounds (Birhane, 2022; Cavazos et al., 2020), leading to potential discrimination in law enforcement and hiring practices.
@@
Real-world consequences exemplify the urgent need to address these challenges at the core of AI development. Ensuring fairness in deep learning models presents complex challenges, primarily due to the black-box nature of these models, which often complicates understanding and interpreting decisions. Moreover, the dynamic and high-dimensional nature of the data involved, combined with nuances in fairness definitions, further complicates the detection and correction of bias. This complexity necessitates the development of more sophisticated, inherently fair algorithms.
@@
Previous mitigation strategies dealing with algorithmic bias – whether through pre-processing, in-processing, or post-processing – have significant limitations. Pre-processing techniques, which attempt to cleanse biased data, are labour-intensive, dependent on expert intervention (Salimi et al., 2019), and only eliminate considered biases. Current in-processing methods frequently lead to unstable models and often rely upon arbitrary definitions of fairness (Caton and Haas, 2020). Post-processing techniques, which adjust model predictions directly, ignore deeper issues without addressing the underlying biases in the data and model. These approaches lack stability, generalisability, and the ability to ensure fairness across multiple metrics (Berk et al., 2017).
@@
In this paper, we introduce FairVIC (Fairness through Variance, Invariance, and Covariance), a novel approach that embeds fairness directly into neural networks by optimising a custom loss function. This function is designed to minimise the correlation between decisions and protected characteristics while maximising overall prediction performance. FairVIC integrates fairness through the concepts of variance, invariance, and covariance during the training process, making it more principled and intuitive, and universally applicable to diverse datasets. Unlike previous methods that often optimise to a chosen fairness metric, FairVIC offers a robust, generalisable solution that introduces an abstract concept of fairness to significantly reduce bias. Our experimental evaluations demonstrate FairVIC’s ability to significantly improve performance in all fairness metrics tested without compromising prediction accuracy. We compare our proposed method against comparable in-processing bias mitigation techniques, such as regularisation and constraint approaches, and highlight the improved, robust performance of our FairVIC model.
@@
  • A novel, generalisable in-processing bias mitigation technique for neural networks.
  • A comprehensive experimental evaluation, using a multitude of comparable methods on a variety of metrics across several datasets, including different modalities such as tabular and text.
@@ 2 Related Work
Pre-processing methods aim to fix the data before training, recognising that bias is primarily an issue with the data itself (Caton and Haas, 2020). In practice, this can be done a number of different ways, such as representative sampling, or re-sampling the data to reflect the full population (Shekhar et al., 2021; Ustun et al., 2019), reweighing the data such that different groups influence the model in a representative way (Calders and Žliobaitė, 2013; Kamiran and Calders, 2012), or generating synthetic data to balance out the representation of each group (Jang et al., 2021). Another set of approaches utilises causal methods to delineate relationships between sensitive attributes and the target variables within the data (Chiappa and Isaac, 2019; Kusner et al., 2017; Russell et al., 2017). Such techniques as these are labour-intensive and do not generalise well, requiring an expert with knowledge of the data to manually process each case of a new dataset (Salimi et al., 2019). They also cannot provide assurances that all bias has been removed – a model may draw upon non-linear/ complex relationships between features that lead to bias, which are hard for the expert/method to spot.
@@
In-processing methods aim to train models to make fairer predictions, even upon biased data. There are a plethora of ways in which this has been done. For example, Celis et al. (2019) and Agarwal et al. (2018) utilise a chosen fairness metric and perform constraint optimisation during training. This has the effect that a single fairness metric needs to be chosen, introducing human bias (Caton and Haas, 2020), and this metric must perfectly capture the bias within the data to effectively mitigate it. Therefore, fairness cannot be achieved across multiple definitions in this way (Caton and Haas, 2020). Another approach involves incorporating an adversarial component during model training that penalises the model if protected characteristics can be predicted from its outputs (Zhang et al., 2018; Wadsworth et al., 2018; Xu et al., 2019). These methods are often effective but their main shortcoming is seen in their instability. Finally, the most relevant comparisons from previous work to our proposed method are regularisation-based techniques that incorporate fairness constraints or penalties directly into the model’s loss function during training. There are a number of ways that this has been done, such as through data augmentation strategies to promote less sensitive decision boundaries (Chuang and Mroueh, 2021) or by incorporating fairness adjustments into the boosting process (Cruz et al., 2023). The performance of these models differs from approach to approach, and those that work by constraining the model by a fairness metric directly suffer from the issue of human bias and misrepresenting the bias within the data/model.
@@
Post-processing techniques involve adjusting model predictions or decision rules after training to ensure fair outcomes. In practice, decision thresholds have been adjusted for different groups to achieve equal outcomes in a particular metric (Hardt et al., 2016). Alternatively, labels near the decision boundary can be altered to favour less biased outcomes (Kamiran et al., 2012, 2018). Calibration (Kim et al., 2018; Noriega-Campero et al., 2019) adjusts the predictions of the model directly so that the proportion of positive instances is equal across each sub-group. This category of methods can oversimplify fairness, and they do not fix the underlying issue within the model. For those techniques that require the specification of a single fairness metric, the same issue applies surrounding this choice as before.
@@
To summarise, there currently lies a number of issues which have not yet been solved in parallel within one technique. These are: stability, generalisability, equal improvements to fairness across metrics (Berk et al., 2017), and built without requirements for user-induced definitions of fairness. In this paper, we solve all these requirements for an effective, generalisable approach to mitigate bias through FairVIC.
@@ 3.1 VICReg
Variance-Invariance-Covariance Regularization (VICReg) (Bardes et al., 2021) has previously been used in self-supervised learning to tackle feature collapse and redundancy. It maximises variance across features to ensure the model produces diverse outputs for different inputs, minimises invariance between augmented representations of the same input to enhance stability, and reduces covariance among features to capture a broader range of information. VICReg is confined to this specific context and objective, and the application of these principles outside of self-supervised methods remains largely unexplored. In contrast, FairVIC extends these principles to supervised learning for bias mitigation. This adaptation addresses the challenges of fairness in decision-making systems, expanding the application of VIC principles beyond their original scope and offering a novel, generalisable solution to fairness in supervised learning models.
@@ 3.2 Group Fairness Metrics
  • Equalized Odds Difference requires that both the True Positive Rate (TPR) and False Positive Rate (FPR) are the same across groups defined by the protected attribute, where $TPR=\frac{TP}{TP+FN}$ and $FPR=\frac{FP}{FP+TN}$ (Hardt et al., 2016). Therefore, we calculate $\max\left(\left|FPR_{u}-FPR_{p}\right|,\left|TPR_{u}-TPR_{p}\right|\right)$, where ${u}$ represents the unprivileged groups and ${p}$ the privileged group and 0 signifies perfect fairness.
  • Average Absolute Odds Difference averages the absolute differences in the false positive rates and true positive rates between groups, defined as $\frac{1}{2}(|FPR_{u}-FPR_{p}|+|TPR_{u}-TPR_{p}|)$, where ${u}$ represents the unprivileged groups and ${p}$ the privileged group, with 0 signifying perfect fairness.
  • Statistical Parity Difference evaluates the difference in the probability of a positive prediction between groups, aiming for 0 to signify perfect fairness. Formally, $DP=|P(\hat{Y}=1|u)-P(\hat{Y}=1|p)|$, where ${u}$ represents the unprivileged groups, ${p}$ the privileged group, and $\hat{Y}=1$ a positive prediction (Dwork et al., 2012).
  • Disparate Impact compares the proportion of positive outcomes for the unprivileged group to that of the privileged group, with a ratio of 1 indicating no disparate impact, and therefore perfect fairness. Denoted as $DI=\frac{P(\hat{Y}=1|u)}{P(\hat{Y}=1|p)}$, where ${u}$ represents the unprivileged groups, ${p}$ the privileged group, and $\hat{Y}=1$ a positive prediction (Feldman et al., 2015).
@@ 3.3 Individual Fairness
While FairVIC aims to increase group fairness, the invariance term promotes direct improvements in individual fairness. This can be observed in our evaluations through counterfactual fairness (Kusner et al., 2017). Counterfactual fairness ensures that decisions made by an algorithm are fair even when considering hypothetical (counterfactual) scenarios. For each individual, the sensitive attribute is switched to assess the model’s ability to perform equally in both the original and counterfactual scenarios.
@@ 4 Approach
We propose FairVIC (Fairness through Variance, Invariance, and Covariance), a novel loss function that enables a model’s ability to learn fairness in a robust manner. FairVIC is comprised of three terms: variance, invariance, and covariance. Optimising for these three terms encourages the model to be stable and consistent across protected characteristics, thereby reducing bias during training. By adopting this broad, generalised approach to defining bias, FairVIC significantly improves performance across a range of fairness metrics. This makes it an effective strategy for reducing bias across various applications, ensuring more equitable outcomes in diverse settings.
@@ 4.1 FairVIC Training
To understand how FairVIC operates, it is crucial to define variance, invariance, and covariance within the context of fairness:
Variance: This term promotes diversity in the latent representations by penalizing low variance in the bottleneck embeddings of the neural network. It ensures the embeddings capture sufficient information, not relying upon a trivial relationship such as the protected characteristic in order to find a solution.
$L_{\text{var}}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\gamma-\sigma(z))$ (1)
where $\sigma(z)$ represents the standard deviation of the embeddings, $\gamma$ is a margin parameter that controls the desired variability, and $N$ is the number of samples.
Invariance: This term ensures the model’s predictions remain consistent when the protected attribute is flipped, promoting individual/counterfactual fairness.
@@
Covariance: This component seeks to reduce the model’s reliance on protected attributes when making predictions, ensuring that decisions are made independently of these attributes. By doing so, it promotes group fairness. The loss function is designed to minimize this covariance, as defined by the following equation:
$L_{\text{cov}}=\frac{\sqrt{\sum_{i=1}^{N}\left(\left(\hat{y}_{i}-\mathbb{E}[\hat{y}]\right)^{\top}\cdot P_{i}\right)^{2}}}{N}$ (3)
where $\hat{y}$ is the model’s prediction, $P$ is the protected attribute, and $N$ is the number of samples.
Algorithm 1 FairVIC Loss Function
1:  Input: Model $M$, Epochs $E$, Batch size $B$, Data $D$, Protected attribute $P$, Weights ($\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}}$)
2:  Output: Trained Model $M$
3:  Initialise $M$
4:  for $e\in\{1,\dots,E\}$ do
5:     Shuffle data $D$
6:     for each batch $\{(X,Y)\}\in D$ with size $B$ do
7:        $\hat{Y}\leftarrow M(X)$
8:        $Z\leftarrow\text{BottleneckLayer}(X)$
9:        Calculate FairVIC Loss:
10:           $L_{\text{acc}}\leftarrow\text{AccuracyLoss}(Y,\hat{Y})$
11:           $L_{\text{var}}\leftarrow\text{VarianceLoss}(Z)$
12:           $L_{\text{inv}}\leftarrow\text{InvarianceLoss}(\hat{Y},M(\text{Flip}(X,P)))$
13:           $L_{\text{cov}}\leftarrow\text{CovarianceLoss}(\hat{Y},P)$
14:           $L_{\text{total}}\leftarrow\lambda_{\text{acc}}L_{\text{acc}}+\lambda_{\text{var}}L_{\text{var}}+\lambda_{\text{inv}}L_{\text{inv}}+$
15:               $\lambda_{\text{cov}}L_{\text{cov}}$
16:        Compute gradients $\nabla L_{\text{total}}\leftarrow\frac{\partial L_{\text{total}}}{\partial M}$
17:        Update model parameters $M\leftarrow M-\alpha\nabla L_{\text{total}}$
18:     end for
19:  end for
During the training of a deep learning model, the model iterates over epochs $E$. Data is shuffled into batches, upon which the model predicts to produce a set of predictions $\hat{Y}$. Typically, the true labels $Y$ and predictions $\hat{Y}$ are then passed into a suitable accuracy loss function (e.g., binary cross-entropy, hinge loss, Huber loss, etc.) and the resulting loss attempts to be minimised by an optimiser.
In the case of FairVIC, in addition to computing a suitable accuracy loss $L_{\text{acc}}$, we also calculate our three novel terms $L_{\text{var}},L_{\text{inv}},\text{and}\;L_{\text{cov}}$ using Equations 12, and 3 respectively. Each of these individual loss terms is then multiplied by its respective weighting factor $\lambda$ and summed to form the total loss $L_{\text{total}}$. Subsequently, gradients are computed, and the optimiser adjusts the model parameters with respect to this combined loss. Further details are provided in Algorithm 1.
The multipliers $\lambda$ enable users to balance the trade-off between fairness and predictive performance, which is typical in bias mitigation techniques. Assigning a higher weight to $\lambda_{\text{acc}}$ directs the model to prioritise accuracy while increasing the weights of $(\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}})$ shifts the focus towards enhancing fairness in the model’s predictions. In our implementation, the lambda coefficients $(\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}})$ are constrained such that their sum equals one. In other words, $\lambda_{\text{acc}}=1-\lambda_{\text{var}}-\lambda_{\text{inv}}-\lambda_{\text{cov}}$. This normalisation ensures the optimisation will not produce multiple solutions in the form $\{k.\lambda_{\text{acc}},k.\lambda_{\text{var}},k.\lambda_{\text{inv}},k.\lambda_{\text{cov}}\},k\in\cal{R}$.

### v3

@@ Abstract
Mitigating bias in automated decision-making systems, particularly in deep learning models, is a critical challenge due to nuanced definitions of fairness, dataset-specific biases, and the inherent trade-off between fairness and accuracy. To address these issues, we introduce FairVIC, an innovative approach that enhances fairness in neural networks by integrating variance, invariance, and covariance terms into the loss function during training. Unlike methods based on predefined criteria, FairVIC abstracts fairness to minimise dependency on protected attributes. We evaluate FairVIC against comparable bias mitigation techniques on benchmark datasets, considering both group and individual fairness, and conduct an ablation study on the accuracy-fairness trade-off. FairVIC demonstrates significant improvements ($\approx 70\%$) in fairness across all tested metrics without significantly compromising accuracy ($\approx-5\%$), thus offering a robust, generalisable solution for fair deep learning across diverse tasks and datasets.
keywords:
[…]
@@ 1 Introduction
With the ever-increasing utilisation of Artificial Intelligence (AI) in everyday applications, neural networks have emerged as pivotal tools for automated decision making systems in sectors such as healthcare [1], finance [2], and recruitment [3]. However, bias in the data–stemming from historical inequalities, imbalanced distributions, or flawed feature representations–are often learned by these models, posing significant challenges to fairness. Such bias can lead to real-world harms. For instance, several studies have shown how bias in facial recognition technologies disproportionately misidentifies individuals of certain ethnic backgrounds [4, 5], leading to potential discrimination in law enforcement and hiring practices.
@@
This highlights the urgent need to address AI bias. Ensuring fairness in deep learning models presents complex challenges, primarily due to the black-box nature of these models, which often complicates understanding and interpreting decisions. Moreover, the dynamic and high-dimensional nature of the data involved, combined with nuances in fairness definitions, further complicates the detection and correction of bias. This complexity necessitates the development of more sophisticated, inherently fair algorithms.
@@
Previous mitigation strategies dealing with algorithmic bias–whether through pre-processing, in-processing, or post-processing–have significant limitations. Pre-processing techniques, which attempt to cleanse biased data, are labour-intensive, dependent on expert intervention [6], and only eliminate considered biases. Current in-processing methods frequently lead to unstable models and often rely upon arbitrary definitions of fairness [7]. Post-processing techniques, which adjust model predictions directly, ignore deeper issues without addressing the underlying biases in the data and model. These approaches lack stability, generalisability, and the ability to ensure fairness across multiple metrics [8].
@@
In this paper, we introduce FairVIC (Fairness through Variance, Invariance, and Covariance), a novel approach that embeds fairness directly into neural networks by optimising a custom loss function. This function is designed to minimise the correlation between binary decisions and binary protected characteristics while maximising overall prediction performance. FairVIC integrates fairness through the concepts of variance, invariance, and covariance during the training process, making it more principled and intuitive, and universally applicable to diverse datasets. Unlike previous methods that often optimise to a chosen fairness metric, FairVIC offers a robust, generalisable solution that introduces an abstract concept of fairness to significantly reduce bias. Our experimental evaluations demonstrate FairVIC’s ability to significantly improve performance in all fairness metrics tested without compromising prediction accuracy. We compare our proposed method against comparable in-processing bias mitigation techniques, such as regularisation and constraint approaches, and highlight the improved, robust performance of our FairVIC model.
@@
  • A novel, generalisable in-processing bias mitigation technique for neural networks;
  • A comprehensive experimental evaluation, using a multitude of comparable methods on a variety of metrics across several datasets, including different modalities such as tabular, text, and image;
@@ 2 Related Work
Pre-processing methods aim to fix the data before training, recognising that bias is primarily an issue with the data itself [7]. In practice, this can be done a number of different ways, such as representative sampling, or re-sampling the data to reflect the full population [9, 10], reweighing the data such that different groups influence the model in a representative way [11, 12], or generating synthetic data to balance out the representation of each group [13]. Another set of approaches utilises causal methods to delineate relationships between sensitive attributes and the target variables within the data [14, 15, 16]. Such techniques as these are labour-intensive and do not generalise well, requiring an expert with knowledge of the data to manually process each case of a new dataset [6]. They also cannot provide assurances that all bias has been removed–a model may draw upon non-linear/ complex relationships between features that lead to bias, which are hard for the expert/method to spot.
@@
In-processing methods aim to train models to make fairer predictions, even upon biased data. There are a plethora of ways in which this has been done. For example, Celis et al. [17] and Agarwal et al. [18] utilise a chosen fairness metric and perform constraint optimisation during training. This requires choosing a fairness metric, introducing human bias and limiting generalisation [7]. Therefore, fairness cannot be achieved across multiple definitions in this way [7]. Another approach involves incorporating an adversarial component during model training that penalises the model if protected characteristics can be predicted from its outputs [19, 20, 21]. These methods are often effective but their main shortcoming is seen in their instability [22]. Finally, the most relevant comparisons from previous work to our proposed method are regularisation-based techniques that incorporate fairness constraints or penalties directly into the model’s loss function during training. There are a number of ways that this has been done, such as through data augmentation strategies to promote less sensitive decision boundaries [23] or by incorporating fairness adjustments into the boosting process [24]. The performance of these models differs from approach to approach, and those that work by constraining the model by a fairness metric directly suffer from the issue of human bias and misrepresenting the bias within the data/model.
@@
Post-processing techniques involve adjusting model predictions or decision rules after training to ensure fair outcomes. In practice, decision thresholds have been adjusted for different groups to achieve equal outcomes in a particular metric [25]. Alternatively, labels near the decision boundary can be altered to favour less biased outcomes [26, 27]. Calibration [28, 29] adjusts the predictions of the model directly so that the proportion of positive instances is equal across each sub-group. These methods oversimplify fairness and ignore model-level bias. For those techniques that require the specification of a single fairness metric, the same issue applies surrounding this choice as before.
@@
To summarise, there currently lies a number of issues which have not yet been solved in parallel within one technique. These are: stability, generalisability, equal improvements to fairness across metrics [8], and built without requirements for user-induced definitions of fairness. In this paper, we solve all these requirements for an effective, generalisable approach to mitigate bias through FairVIC.
@@ 3.1 VICReg
Variance-Invariance-Covariance Regularization (VICReg) [30] has previously been used in self-supervised learning to tackle feature collapse and redundancy. It maximises variance across features to ensure the model produces diverse outputs for different inputs, minimises invariance between augmented representations of the same input to enhance stability, and reduces covariance among features to capture a broader range of information. VICReg has mostly been confined to self-supervised learning, with little exploration beyond. To this extent, FairVIC reworks the VIC components completely to serve its application in supervised learning for bias mitigation, a principle that has remained so far unexplored. This adaptation addresses the challenges of fairness in decision-making systems, expanding the application of VIC principles beyond their original scope and offering a novel, generalisable solution to fairness in supervised learning models.
@@ 3.2 Group Fairness Metrics
  • Equalized Odds Difference requires that both the True Positive Rate (TPR) and False Positive Rate (FPR) are the same across groups defined by the protected attribute, where $TPR=\frac{TP}{TP+FN}$ and $FPR=\frac{FP}{FP+TN}$ [25]. Therefore, we calculate $\max\left(\left|FPR_{u}-FPR_{p}\right|,\left|TPR_{u}-TPR_{p}\right|\right)$, where u{u} represents the unprivileged groups and p{p} the privileged group and 0 signifies perfect fairness.
  • Average Absolute Odds Difference averages the absolute differences in the false positive rates and true positive rates between groups, defined as $\frac{1}{2}(|FPR_{u}-FPR_{p}|+|TPR_{u}-TPR_{p}|)$, where u{u} represents the unprivileged groups and p{p} the privileged group, with 0 signifying perfect fairness.
  • Statistical Parity Difference evaluates the difference in the probability of a positive prediction between groups, aiming for 0 to signify perfect fairness. Formally, $DP=|P(\hat{Y}=1|u)-P(\hat{Y}=1|p)|$, where u{u} represents the unprivileged groups, p{p}–the privileged group, and $\hat{Y}=1$–a positive prediction [31].
  • Disparate Impact compares the proportion of positive outcomes for the unprivileged group to that of the privileged group, with a ratio of 1 indicating no disparate impact, and therefore perfect fairness. Denoted as $DI=\frac{P(\hat{Y}=1|u)}{P(\hat{Y}=1|p)}$, where u{u} represents the unprivileged groups, p{p}–the privileged group, and $\hat{Y}=1$–a positive prediction [32].
@@ 3.3 Individual Fairness
While FairVIC aims to increase group fairness, the invariance term promotes direct improvements in individual fairness. This can be observed in our evaluations through counterfactual fairness [15]. Counterfactual fairness ensures that decisions made by an algorithm are fair even when considering hypothetical (counterfactual) scenarios. For each individual, the sensitive attribute is switched to assess the model’s ability to perform equally in both the original and counterfactual scenarios.
@@ 4 Approach
We propose FairVIC (Fairness through Variance, Invariance, and Covariance), a novel loss function that enables a model’s ability to learn fairness in a robust manner. FairVIC is comprised of three terms: variance, invariance, and covariance. Optimising for these three terms encourages the model to be stable and consistent across protected characteristics, thereby reducing bias during training. This broad, generalised approach to defining bias improves performance across a range of fairness metrics. This makes it an effective strategy for reducing bias across various applications, ensuring more equitable outcomes in diverse settings.
@@ 4.1 FairVIC Training
To understand how FairVIC operates, it is crucial to define variance, invariance, and covariance in their adapted forms: each is based on a classical statistical concept but used to express a specific fairness objective within the model. We consider binary classification tasks, where the model output $\hat{y}\in[0,1]$ represents the predicted probability of the favourable class, obtained via a sigmoid activation.
Variance: This term promotes diversity in the latent representations by penalising low variance across features in the bottleneck embeddings, denoted $z\in\mathbb{R}^{d}$, where $z$ is the output of the encoder for each input. It ensures the embeddings capture sufficient information, rather than collapsing to a trivial relationship such as the protected characteristic.
$L_{\text{var}}=\frac{1}{d}\sum_{j=1}^{d}\max(0,\gamma-\sigma_{j}(z))$ (1)
where $\sigma_{j}(z)$ represents the standard deviation of the $j$-th feature across the batch, we set $\gamma$ to 1.0 as a margin parameter to encourage unit-scale variability in the latent representation, and $d$ is the number of features in the embedding.
Invariance: This term ensures the model’s predictions remain consistent when the protected attribute is flipped, promoting individual (counterfactual) fairness.
@@
Covariance: This component seeks to penalise any contribution of the protected attribute to the decision process of the classifier, ensuring that predictions are not systematically skewed for the privileged group. By doing so, it promotes group fairness. The loss function is designed to minimise this covariance, as defined by the following equation:
$\color[rgb]{0,0,0}L_{\text{cov}}=\frac{1}{N}\sqrt{\sum_{i=1}^{N}\left(\left(\hat{y}_{i}-\mathbb{E}[\hat{y}]\right)\cdot P_{i}\right)^{2}}$} (3)
where $\hat{y}\in(0,1)$ is the model’s softmax output, $P_{i}\in\{0,1\}$ is the binary protected attribute with $P_{i}=1$ denoting the privileged group, and $N$ is the number of samples. In general, $\mathbb{E}[\hat{y}$] reflects the empirical average of predictions across the batch, though it may approach 0.5 in a balanced and well-calibrated setting. During training, this loss reduces the deviation of privileged group predictions from the batch mean. In practice, it tends to soften overly confident predictions for the privileged group without necessarily changing the predicted class. This term does not directly affect accuracy, which is controlled by a separate loss, but its interaction with the accuracy objective discourages reliance on group membership in decision-making, thereby promoting group fairness.
Together, alongside a suitable accuracy loss, FairVIC jointly optimises for the total loss equation seen in Equation 4.
$\color[rgb]{0,0,0}L_{\text{total}}=\lambda_{\text{acc}}L_{\text{acc}}+\lambda_{\text{var}}L_{\text{var}}+\lambda_{\text{inv}}L_{\text{inv}}+\lambda_{\text{cov}}L_{\text{cov}}$} (4)
where each $\lambda$ is a non-negative weighting coefficient balancing the contribution of its corresponding term, subject to the normalisation constraint $\lambda_{\text{acc}}+\lambda_{\text{var}}+\lambda_{\text{inv}}+\lambda_{\text{cov}}=1$.